Express the confidence interval 78.9 % < p < 94.9 % 78.9 % < p < 94.9 % in the form of ˆ p ± M E p ^ ± M E . % ± ± %
Answer #1

Values are given as,

Lower limit = 78.9%

Upper limit = 94.9%

Calculating sample proportion:

\hat{p}= (upper limit +lower limit)/2

= (94.9+78.9)/2

= 173.8/2

= 86.9%

Calculating margin of error :

E = (upper limit - lower limit)/2

= (94.9-78.9)/2

= 16/2

= 8%

Calculating confidence interval :

CI= \hat{p}\pm ME
CI= 86.9% \pm 8%

Required confidence interval : 86.9% \pm 8%